which number is closer to 1: 9/10 or 10/9?
Answer by stanbon(75887)

which number is closer to 1: 9/10 or 10/9?
-------------------
1-(9/10) = 1/10
--------------------
1-(10/9) = (9/9)-(10/9) = -1/9
-----
=====
1/10 is smaller then 1/9 so (9/10) is closer
=============================================
Cheers,
Stan H.
====================